A round pizza, with a diameter of 14 inches, is cut into 12 equal slices. A square pizza, with side length 12 inches, is cut into 9 equal squares. Which pizza slice, sector or square,has the greatest area? How much greater is it?
Answer by texttutoring(324)

First, find the area of the round pizza:

A=Pi*r^2
A=Pi*(7)^2
A=49*Pi

The area of a single slice would be 1/12 of the whole area:
Aslice = 49*Pi/12 = 12.83 inches squared


Now find the area of the square pizza:

A=x^2
A=12^2
A=144

We are cutting it into 9 pieces, so to find the area of one piece:
Aslice = 144/9
Aslice = 16 inches squared

So the square slice is larger, by (16-12.83) 3.17 square inches.